# SAP-CoPE: Social-Aware Planning using Cooperative Pose Estimation with Infrastructure Sensor Nodes

### v1

@@ VII-A2 Results
Camera Pose 0.1166 / 0.1308 8.4912 / 10.2322
Camera-Lidar Pose 0.0555 / 0.0640 8.3785 / 10.0688
VII-A3 Error Analysis
Position Error
The integration of LiDAR data significantly reduces position estimation errors, as measured by MAE and RMSE, by 52.40% and 51.07% Specifically, the Camera-LiDAR method (MAE = 0.0555m, RMSE = 0.0640m) outperforms the camera-only approach (MAE = 0.1166m, RMSE = 0.1308m). This improvement is evident in the spatial distribution of errors (Fig. 6), where the Camera-LiDAR method demonstrates consistently lower positional errors throughout the evaluated region. Nonetheless, the camera-only method still achieves noteworthy performance, underscoring its effectiveness even without LiDAR data.
Yaw Error
Yaw errors are relatively consistent across both methodologies, with the Camera-LiDAR fusion method showing only marginally lower errors (MAE = 8.3785) compared to the camera-only method (MAE = 8.4912). This suggests that orientation estimation benefits minimally from the addition of LiDAR data, in contrast to the substantial improvements observed in positional accuracy. Despite this minor difference, the camera-only yaw estimation remains robust and effective for practical use.
Figure 6: Spatial distribution of position errors. (a) Camera-only pose estimation. (b) Camera-Lidar fusion.
Figure 7: Spatial distribution of yaw errors. (a) Camera-only pose estimation. (b) Camera-Lidar fusion.
VII-A4 Discussion
The results indicate that integrating LiDAR data significantly improves position accuracy but has a limited effect on yaw estimation. This is likely because LiDAR provides precise depth information, which aidis in localization, whereas yaw estimation largely depends on 2D visual cues.
VII-B Overall Experiments
The proposed SAP-CoPE framework was evaluated through three scenarios: obstacle and human interaction, yielding, and group interaction. Figure 8 illustrates an example of the experimental setup, including RViz visualization and left and right images from the ISN. A detailed visualization, including simulation results and all experiments, is available at https://github.com/HopeYless/SAP-CoPE-Project.
Figure 8: Instance in a real-life experiment on the robot, showing the RViz visualization and images captured by an ISN
VII-B1 Scenario One: Obstacle and Human Interaction
Figure 9: Experiment Results of Obstacle and Human Interaction Scenario
In this scenario, the robot encounters two obstacles obstructing its reference path. The first obstacle, representing a stationary object, is positioned at the top right as a green rectangle. The second obstacle is a human, moving from left to right along the scene, represented by a green circle, with the black line indicating their heading angle. The robot’s reference trajectory is defined as the centerline of the hallway, with a reference heading angle of constant 180 degrees. The reference speed for both wheels is set to a constant 0.8 m/s.
In Fig. 9 (a), the solid black markers (circles, triangles, diamonds, and squares) represent different time stamps of the robot and the human. The robot is shown in a blue rectangle. The plot demonstrates that the robot successfully avoided both obstacles, maintaining minimal deviation from its reference trajectory. Fig. 9 (b) and (c) present the robot’s heading angle and wheel linear velocity, respectively. The heading angle plot confirms that the robot consistently tracked the reference signal throughout the experiment, while the velocity plot shows that the robot follows to the reference wheel velocity.
Figs. 9 (d) and (e) demonstrated the optimal acceleration and steering angle generated by the proposed MPC controller. All values remained within the defined constraints, confirming the effectiveness of the control strategy.
VII-B2 Scenario Two: Yielding
Figure 10: Experiment Results of Yielding Scenario
In the second scenario, the planning algorithm was tested to determine whether the robot could appropriately yield to other humans and reduce speed when there was not enough clearance to pass. This scenario involved two humans: one stationary, facing to the right, and another approaching the robot rapidly from behind. The reference trajectory was defined the same as in the previous case.
Fig. 10 (a) demonstrates the robot’s ability to accurately identify the situation. Initially, the robot moved downward to yield to the fast-approaching individual. At the 18th second, both the stationary and dynamic individuals occupied positions that left not enough space for the robot to proceed. In response, the robot correctly decided to reduce speed at an appropriate distance in front of the stationary individual and wait for the dynamic person to pass. Once the space cleared, the robot continued to track its defined reference motion.
Fig. 10 (b) and (c) illustrate the heading angle and wheel linear velocity of the robot, respectively. Both plots show the robot closely following the defined reference signal. Lastly, Fig. 10 (d) and (e) show the optimal acceleration and steering angle generated by the proposed MPC, all of which fall within defined constraints.
VII-B3 Scenario Three: Group Interaction
In the third scenario, the planning algorithm was tested to verify if the algorithm is able to identify a group of people and output appropriate actions. In this scenario, two people were initially stationary and facing each other, leaving sufficient clearance for the robot to pass. The reference trajectory was defined the same as in the previous cases.
Fig. 11 (a) demonstrates the robot’s ability to identify the group of people. Initially, the robot came to a complete stop in front of the two people, even though the clearance between them was enough for the robot to pass. However, the robot recognized that passing between the two individuals could disturb both of them. Once the person in front had walked away, the robot resumed its operation and followed the defined reference trajectory.
Fig. 11 (b) and (c) illustrate the heading angle and wheel linear velocity of the robot, respectively. Both plots show the robot closely following the defined reference signal. Note, in Fig. 11 (c), the plot demonstrates the robot comes to a complete stop around the 7th second. Lastly, Fig. 11 (d) and (e) show the optimal acceleration and steering angle generated by the proposed MPC, all of which fall within defined constraints.
Figure 11: Experiment Results of Group Interaction Scenario
VIII Conclusions and Future Work
@@
In this paper, we propose a SAP-CoPE framework for autonomous driving systems operating in human-populated environments. By leveraging a cooperative infrastructure sensor network and a novel 3D human pose estimation method, our approach effectively addresses perception limitations and occlusion challenges of accurate human intention recognition and safe navigation. Additionally, the integration of human-pose-based PS concepts into an MPC controller enables the generation of trajectories that prioritize both safety and human comfort. Extensive evaluations in both simulation and real-world environments validate the effectiveness of our proposed SAP-CoPE framework by producing socially aware trajectories. Future work will focus on enhancing system robustness in highly dynamic environments and further improving human intention prediction with the combination of learning-based algorithms.

### v2

@@ VI-A1 Evaluation Metrics
3DNet 0.1795 / 0.2163 8.5629 / 12.1001
Metrabs 0.2294 / 0.2926 8.1819 / 11.2037
Cam-only 0.1098 / 0.1308 8.5199 / 11.7416
Cam+LiDAR 0.0892 / 0.1090 7.9168 / 11.0136
Cam-only (coop) 0.1087 / 0.1237 8.1727 / 10.2264
Cam+LiDAR (coop) 0.0771 / 0.0862 6.8607 / 8.9542
TABLE III: Delay mitigation evaluation results
Delay Mode F1 Recall Avg. DE Avg. AE
0ms Baseline 0.9456 0.9512 0.0829 7.2475
Delay-Aware 0.9485 0.9739 0.0818 5.9557
50ms Baseline 0.9456 0.9510 0.0864 7.2314
Delay-Aware 0.9480 0.9739 0.0840 6.1045
100ms Baseline 0.9455 0.9510 0.0975 7.2110
Delay-Aware 0.9480 0.9738 0.0876 6.2546
200ms Baseline 0.9449 0.9513 0.1439 7.2365
Delay-Aware 0.9479 0.9740 0.0993 6.5910
VI-A2 Results
We evaluate both pose estimation accuracy and system robustness under computation and communication delay. For pose estimation, we compare six configurations: learning-based baselines (3DNet[20], Metrabs[29]), single-node camera-only and camera–LiDAR fusion, and their cooperative multi-node counterparts. Quantitative metrics are shown in Table II, and spatial error distributions are visualized in Figs. 5 and 6.
Position error: Integrating LiDAR measurements consistently improves localization accuracy over camera-only methods. Single-node camera–LiDAR fusion reduces MAE from 0.1098 m to 0.0892 m. When cooperative perception is used to refine the pose using multi-view data, performance further improves: Cam+LiDAR (coop) achieves the lowest error of 0.0771 m MAE and 0.0862 m RMSE, outperforming all baselines. The advantages of the proposed method can be further highlighted by the spatial heatmaps (Fig. 5). The learning-based methods show significant degradation in distant areas, mainly because they generalize poorly under highly varying perspective views. In contrast, the proposed methods, even the camera-only version, maintains consistently lower errors. This improvement comes from our uncertainty-aware pose estimation, which explicitly models the camera projection process and propagates measurement uncertainty through the geometry, yielding more reliable estimates in distant and low-observability regions.
Yaw error: Orientation estimation shows smaller but consistent improvements. While LiDAR provides limited direct orientation cues, cooperative multi-view observations help stabilize heading estimation. Cam+LiDAR (coop) reduces yaw MAE to 6.86, compared to 8.52 for camera-only and over 8 for learning-based baselines. This demonstrates that multi-view geometric consistency is particularly beneficial for resolving ambiguous orientations under occlusions.
VI-A3 Delay Robustness Evaluation
In our system, local perception requires an average of 42 ms, while communication latency is approximately 3 ms over local Wi-Fi and 53 ms over 5G; global fusion and tracking add a further 2 ms [23]. Consequently, practical deployments may experience total latency ranging from tens to hundreds of milliseconds. To quantify robustness to latency, we evaluate performance under simulated communication delays from 0 to 200 ms, with results summarized in Table III.
The baseline method directly fuses all received observations without multi-object tracking and delay compensation, while the proposed delay-aware pipeline incorporates timestamp-based delay correction and motion prediction. Across all delay settings, the delay-aware pipeline preserves stable perception quality and consistently outperforms the baseline, yielding higher recall as well as lower displacement error (Avg. DE) and angular error (Avg. AE). The advantage becomes more obvious as delays increase: at 200 ms, the baseline position error rises to 0.1439 m, whereas the delay-aware approach limits it to 0.0993 m. Similarly, the yaw error decreases from 7.24 to 6.59. These results indicate that timestamp-based compensation coupled with motion prediction effectively mitigates stale observations and improves temporal consistency under significant communication latency.
VI-A4 Discussion
The experiments highlight:
  • LiDAR fusion significantly improves positional accuracy by providing reliable depth constraints.
  • Cooperative multi-node perception reduces occlusion and improves both position and heading estimation through multi-view redundancy.
  • The delay-aware fusion strategy maintains high accuracy under latency, crucial for real-world deployments with computation and communication delays.
Together, these components enable accurate and robust human pose estimation suitable for real-time indoor robot navigation in crowded and communication-constrained environments.
VI-B Combined Experiments
As the performance of the perception of the SAP-CoPE system was already validated in the previous section, the main focus in this section is focus on the motion planning system of the proposed system by using the proposed perception system. The perception results with computation and communication delay mitigation are sent to all motion planners, including the baseline planners. The experiments were conducted in a hallway-like environment with a straight and a 90-degree turn setup. During the experiments, the robot should follow the predefined reference trajectory while proactively avoiding humans.
VI-B1 Mobile Robot
The robot is retrofitted based on an existing medical bed. A picture of the robot is shown in Fig. 2, in the ”Application Layer” block. To ensure optimal traction and handling, the robot incorporates a two-wheel-drive system (one wheel in the front and one in the back) and an independent steering system, allowing each wheel to turn independently for omnidirectional mobility.
VI-B2 Experiments Setup
Figure 7: Experiment setup with all participants. Images captured from a) ISN1’s left camera. b) ISN1’s right camera. c) ISN2’s left camera. d) ISN2’s right camera.
TABLE IV: Robot and MPC Parameters. Units for velocity (or acceleration) are in [m/s] (or [m/s2]), and steering angles (or steering rate)are in [rad] (or [rad/s]).
Symbol Value Symbol Value
$l_{f},l_{r}$ 1.2 m $u_{c\max}$ $\left[0.7,0.7,\frac{\pi}{2},\frac{\pi}{2}\right]$
$T_{s}$ 0.1 sec $u_{c\min}$ -$\left[0.7,0.7,\frac{\pi}{2},\frac{\pi}{2}\right]$
$N_{p}$ 2020 $\Delta u_{c\max}$ $\left[1,1,\frac{\pi}{24},\frac{\pi}{24}\right]$
$N_{c}$ 10 $\Delta u_{c\min}$ -$\left[1,1,\frac{\pi}{24},\frac{\pi}{24}\right]$
TABLE V: Statistics Results Comparison between the Proposed Method and Other Methods
. Clearance (Avg. / Std.) [m] Traveled Distance (Avg. / Std.) [m] Traveled Time (Avg. / Std.) [sec.] # Human SFM MPC + PF Proposed SFM MPC + PF Proposed SFM MPC + PF Proposed 2 0.23 / 0.055 0.39 / 0.052 0.37 / 0.063 12.04 / 0.18 12.19 / 0.15 12.18 / 0.19 18.10 / 0.22 17.78 / 0.15 17.75 / 0.18 4 0.089 / 0.051 0.21 / 0.075 0.27 / 0.064 12.08 / 0.12 12.22 / 0.23 12.37 / 0.19 18.61 / 0.42 18.93 / 0.74 18.35 / 0.28 6 0.020 / 0.028 0.11 / 0.087 0.26 / 0.042 21.16 / 0.62 21.86 / 0.59 20.76 / 0.18 34.18 / 2.82 33.43 / 1.58 30.73 / 0.39
The proposed perception system and planner were implemented and tested in experiments using the C++ and Python platforms. To verify our proposed planner, we directly send the generated control inputs to the robot for simplicity. The parameters for the robot and the MPC motion planner are summarized in the table IV. All computations in the CL, including global perception fusion and motion planning algorithm are carried out on a Linux-based desktop equipped with an Intel i7-14700F CPU and 64GB of RAM.
The participant’s initial position (numbered blue circles), movement direction (yellow arrow), and ISNs’ location can be found in Fig. 7. Three experimental scenarios were set up to validate the proposed system. In scenario (1), which involved 2 participants, the robot was asked to follow the reference path and stop at ”Goal 1”, while participants 3 and 4 were involved. In scenario (2), which involved 4 participants, the goal of the robot and the movement of participants 3 and 4 remained the same, but with the addition of participants 2 and 4. In scenario (3), which involved all participants, the goal of the robot was set to ”goal 2”, and participants 1 to 4 remained the same, but with the addition of participants 5 and 6.
VI-B3 Evaluation Metrics
Three evaluation metrics were used to validate the performance of the proposed motion planning algorithm when compared with other baselines. a) Clearance: this metric measures the minimum distance between the robot and the human (between borders). Lower values indicate the robot is close to the human, and may increase the psychological pressure on the human and collision risk. b) Travel Distance: this metric measures the total travelled distance of the robot from the starting position to the goal, which reflects the travel efficiency of the robot. c) Travel Time: this metric measures the total travelled time of the robot from the start position to the goal, which also reflects the motion smoothness. A short travel time indicates the robot can maintain a defined speed without too much deceleration or stopping.
VI-B4 Results
(a) SFM
(b) Proposed Method
Figure 8: 2-Participant Experiment Scenario. Trajectory plots for a) Social Force Model (SFM); b) Proposed method that combines MPC and PS model
Figure 9: Proposed Motion Planning Method Computation Time Distribution
The proposed planner method, among other selected baseline methods, was tested for each scenario described in Fig. 7. The first baseline is the conventional Social Force Model (SFM) [12]. Some modifications were made to allow the algorithm to consider the shape of the robot as a rectangle. The second baseline is the MPC with the Potential Field (MPC + PF) [26]. Each method for each scenario was repeated 10 times to reduce the effect of randomness.
The results in Table V demonstrate that the proposed method achieved best overall performance than the other methods in general; specifically, the proposed method maintained an appropriate clearance to participants, a reasonable total travelled distance and the shortest travelled time through all experiments. The results also show that the clearance is decreasing for the baseline methods as the number of participants increases, but the proposed method can maintain a proper clearance. More detailed discussion of the results can be found in the later section.
On the other hand, Fig. 9 presents an analysis of the proposed planner’s computational performance across three experimental scenarios with varying numbers of dynamic obstacles. The visualization employs violin plots (mirrored distribution plot), combined with box plot elements, to comprehensively represent the distribution characteristics of computation times.
By observing the figure, one can see that the 2-Human experiment exhibits the tightest distribution (median 11 ms, mean 11.5 ms) with minimal variability, while the 4-Human scenario shows a moderate increase (median 14 ms, mean 15 ms) with occasional outliers up to 36 ms, indicating periodic solver convergence challenges. The 6-Human experiment demonstrates the highest computational burden (median 19 ms, mean 18.7 ms) with broader distribution spanning 13-26 ms and outliers approaching 37 ms, reflecting the increased optimization complexity of dense multi-obstacle configurations where additional solver iterations are occasionally required. In summary, all scenarios maintain computation times (including perception and motion planner) well below the MPC loop periods (100 ms in our case) for autonomous navigation systems, confirming real-time feasibility.
VI-B5 Discussion
As shown in Table V, the conventional SFM algorithm results in the smallest clearance among all compared methods. This is because the equivalent steering angles and wheel velocities of the front and rear wheels are computed based on the optimal CG. velocity and heading angle obtained from the algorithm; consequently, the robot’s omnidirectional maneuverability is not explicitly considered. As illustrated in Fig. 8, due to the robot’s length, participants may need to actively dodge the robot when the SFM is applied. In contrast, when omnidirectional motion capability is incorporated, the robot can smoothly avoid all participants without causing any disturbances.
Furthermore, it is worth noting that the MPC+PF method exhibits relatively long travel times in the 4-human and 6-human scenarios. This is because the potential field employed in this approach is more conservative than that of the PS field; consequently, the robot occasionally needs to reduce its speed when navigating through crowds. This behaviour is also reflected in the reduction of the clearance metric, particularly in the 6-participant scenarios. Although this characteristic allows the robot to maintain larger distances from participants in relatively sparse environments, it may pose safety concerns in densely populated settings, as unexpected deceleration could cause participants to collide with the robot, especially for those approaching the robot from behind.
In summary, the SFM has relatively worse performance because it does not account for the robot’s omnidirectional maneuverability. The MPC+PF method performs comparably to the proposed approach in less crowded scenarios and can even maintain larger clearances; however, due to its symmetric field representation, its performance decreases significantly in more challenging and densely populated environments.
VII Conclusions and Future Work
@@
In this paper, we propose a SAP-CoPE framework for autonomous driving systems operating in human-populated environments. By leveraging a cooperative infrastructure sensor network and a novel 3D human pose estimation method, our approach effectively addresses perception limitations and occlusion challenges of accurate human intention recognition and safe navigation. Additionally, the integration of human-pose-based PS field into an MPC controller enables the generation of trajectories that prioritize human comfort. Extensive evaluations in real-world environments validate the effectiveness of our proposed SAP-CoPE framework. Due to the modular design and real-time performance, our proposed SAP-CoPE framework has strong potential for deployment in practical applications such as indoor service robots, healthcare logistics, and autonomous navigation in shared public spaces. Future work will focus on enhancing system robustness in highly dynamic environments and further improving human intention prediction with the combination of learning-based algorithms.